Summary: Overall a productive run; problems with camera and telescope jump sometimes frustrating; but weather was beautiful and a fair amount of data taken.

Team: Xiaohui Fan, Linhua Jiang, Fuyan Bian (first half) Ben Wilner, Michael Cooper (second half) helped by Olga Kuhn (supporting astronomer) and John Hill (from Tucson)

Weather: this run has the best weather condition of all runs we have so far. We were able to open the dome for all six nights; although most nights were not photometric, only in one half night did the cloud cover become serious enough to affect data quality. For majority of the time, we had 0.6 - 1.0" seeings.

Camera: this run was highlighted by a series of, sometimes serious, camera issues, mostly has something to do with rotator (later identified as a problem with a flipper) and the science PC for the red. We lost about a total of about 5 hours due to camera issues. Guiding performance similar as before with a fraction of images elongated (in addition to the telescope jumps).

Telescope: it behalved most of the time. The main significant issue is that the telescope elevation mount jumps (by ~1-2") very often, resulting in about 1/5-6 images trailed or have multiple peaks. This is a major loss of efficiency.

OBs: Not counting the ones that were seriously affected by clouds, we did:
night 1: 3 EGS, 1 Bootes, 2 Herc
2: 2 EGS, 4 Bootes, 3 Herc
3: 3 EGS, 5 Bootes, 4 Herc
4: 4 EGS, 4 Bootes, 3 Herc
5: 4 Herc
6: 2 EGS, 4 Bootes, 5 Herc

so the total exposure time is: 19.4 hours, out of about 34 hours of available time (excluding camera down time and weather loss).

Note some of the Herc fields have to be repeated with blue only to ensure guiding.

Summary of AZ programs over the three runs:

Bootes: we have completed AZ share of the Bootes survey (2 nights). All data are being reduced and first spectroscopic follow-up ongoing.

Dwarf Galaxies: reasonably good data collected in both dwarf galaxies. UMaII mostly done, Herc half down.

NGC galaxies: program completed.

Extended Groth Strip: about 1/3 of data collected.

A few SN OBs were observed during AZ runs as well.

As always, many thanks to the LBC and LBTO staff. We will be back in the Fall!

Xiaohui Fan for all observers

-- XiaohuiFan - 03 Jun 2008
Topic revision: r1 - 03 Jun 2008, XiaohuiFan
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback